If the sample contains a combined total of 1.70 mol and has a total mass of 64.2 g, what are the percentages of CO2 and Ne in the sample by mass?
Answer #1

Let number of mol of CO2 be x

Then mol of Ne = 1.70-x

Molar mass of CO2,

MM = 1*MM(C) + 2*MM(O)

= 1*12.01 + 2*16.0

= 44.01 g/mol

Molar mass of Ne = 20.18 g/mol

Total mass = x*44.01 + (1.70-x)*20.18 = 64.2

x*44.01 + (1.70-x)*20.18 = 64.2

44.01*x + 34.31 - 20.18*x = 64.2

23.82*x = 29.89

x = 1.25

Mass of CO2 = 1.25*44.01 = 55.0 g

Mass % of CO2 = mass of CO2 * 100 / total mass

= 55.0 * 100 / 64.2

= 85.7 %

Mass % of Ne = 100 - 85.7 = 14.3 %

Answer: 85.7 %, 14.3 %